# Palu: Compressing KV-Cache with Low-Rank Projection

### v1

@@ Abstract
KV-Cache compression methods generally sample a KV-Cache of effectual tokens or quantize it into lower bits. However, these methods cannot exploit the redundancy of the hidden dimension of KV tensors. This paper investigates a unique hidden dimension approach called Palu, a novel KV-Cache compression framework that utilizes low-rank projection. Palu decomposes the linear layers into low-rank matrices, caches the smaller intermediate states, and reconstructs the full keys and values on the fly. To improve accuracy, compression rate, and efficiency, Palu further encompasses (1) a medium-grained low-rank decomposition scheme, (2) an efficient rank search algorithm, (3) a low-rank-aware quantization algorithm, and (4) matrix fusion with optimized GPU kernels. Our extensive experiments with popular LLMs show that Palu can compress KV-Cache by more than 91.25% while maintaining a significantly better accuracy (up to 1.19 lower perplexity) than state-of-the-art KV-Cache quantization methods at a similar or even higher memory usage. When compressing KV-Cache for 50%, Palu delivers up to 1.61$\times$ end-to-end speedup for the attention module. Our code is publicly available at: https://github.com/shadowpa0327/Palu.
Palu: Compressing KV-Cache with Low-Rank Projection
[…]
footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Large language models (LLMs) are revolutionizing the AI industry and providing a high-level intelligence that previous machine learning (ML) models could not achieve. To speed inference, caching key-value states (KV-Cache) in memory is a simple yet effective technique. However, the size of the KV-Cache can grow rapidly, straining memory capacity and bandwidth; further, the memory-bounded characteristics of the decoding stage also limit inference speed when loading KV-Cache data. Therefore, KV-Cache compression has become a central research topic for building efficient LLMs.
Figure 1: Palu’s low-rank projection method for KV-Cache reduction. A weight matrix W of linear projection is decomposed into two low-rank matrices. Input X is down-projected to latent representation H, which is cached. Y is reconstructed from H on-the-fly using the up-projection matrix B.
Research on KV-Cache compression falls into two categories: quantization and token eviction. Quantization methods aim to use less bit-width to represent each datum; token eviction techniques seek ways to retain only a partial set of KV-Cache. However, both categories fail to explore the hidden dimensions of KV-Cache where high redundancy often occurs.
This work examines the problem from a fundamentally different perspective, i.e., ways to reduce the hidden dimensions. We introduce Palu, a KV-Cache compression framework that reduces the size of the KV-Cache through low-rank projection. Although low-rank projection has been shown to improve LLM efficiency, e.g., by reducing model size (Yuan et al., 2023; Wang et al., 2024) or minimizing memory footprint for finetuning (Hu et al., 2022; Dettmers et al., 2023), it has not been studied for compressing the KV-Cache, especially in post-training.
@@
A naive way to compress the KV-Cache with low-rank projection is by directly mapping cached matrices into low-rank space (Jolliffe and Cadima, 2016; Zhao et al., 2024). However, this approach imposes an unacceptably heavy overhead of computing the decomposition matrices during runtime, after the prefill stage. To avoid this, Palu statically decomposes the key and value-projection weight matrices and caches the latent representations of the low-rank decomposition (Fig. 1), instead of directly decomposing the KV-Cache. This innovative design enables Palu to reduce memory while substantially mitigating runtime overhead.
Despite its promising memory reduction, we found that naively decomposing the key and value projection matrices along the head dimension causes a substantial loss in accuracy. Although we found that decomposing the entire matrix at one time can maintain accuracy, it significantly increases reconstruction overhead. To find a balance, Palu proposes a medium-grained, group-head low-rank decomposition optimization that is both accurate and less costly in terms of reconstruction.
@@
For LLMs, each linear projection module has a different sensitivity to compression Sharma et al. (2023); Yuan et al. (2023). To exploit the sensitivity and improve accuracy, we design an efficient rank search algorithm based on Fisher information. Our algorithm automatically assigns a larger rank for important matrices and smaller ranks for less critical ones, boosting accuracy at the same compression rate.
In addition to low-rank decomposition, we integrate quantization into Palu. We found that low-rank decomposition often induces severe outliers in the latent representation. Inspired by recent LLM quantization methods Tseng et al. (2024); Ashkboos et al. (2024b), we use Hadamard transformation matrices with a low-rank-aware quantization algorithm to eliminate outliers and increase quantization accuracy. To avoid runtime overhead, Palu fuses the transformation matrices into the forward and backward decomposed matrices, concealing the overhead completely.
To enhance inference latency, we integrate the decomposed matrix into the subsequent linear layer, thereby mitigating the online reconstruction overhead. In scenarios where RoPE is applied and matrix fusion is infeasible, we developed an optimized GPU kernel to accelerate the reconstruction.
@@
We evaluate Palu on popular LLMs and benchmarks. Our experiments show that Palu maintains strong zero-shot accuracy and perplexity with up to 50% low-rank compression. Combining low-rank compression and quantization, Palu compresses KV-Cache by over 91.25% (11.4$\times$ compression) while maintaining perplexity. Compared to a state-of-the-art KV-Cache quantization (Hooper et al., 2024) with 87.5% compression, Palu achieves a significantly lower perplexity of 1.19. Additionally, with 50% compression without quantization, Palu demonstrates a 1.61$\times$ end-to-end speedup, highlighting Palu’s ability to reduce memory usage and enhance speed for LLMs.
@@
  • Palu, a new post-training KV-Cache compression framework that caches low-rank latent representations.
  • Group-head low-rank decomposition (G-LRD), an optimization for balancing accuracy and efficiency.
@@
  • A low-rank-aware quantization algorithm that eliminates low-rank-induced outliers and imposes zero runtime overhead.
@@ 2.2 Singular Value Decomposition (SVD)
SVD (Jolliffe and Cadima, 2016) is a commonly used technique for computing the low-rank approximation for a given matrix. We now introduce our specific use of SVD for Palu’s default low-rank decomposition method.
Given a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD decomposes $\mathbf{W}$ into three matrices: $\mathbf{W}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}$. Here, $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices containing the left and right singular vectors, respectively. The matrix $\boldsymbol{\Sigma}$ is a diagonal matrix that consists of singular values. We describe the decomposition as
$\mathbf{W}\approx\mathbf{A}\mathbf{B},\hskip 5.0pt\mathbf{A}=\mathbf{U}_{r}\sqrt{\boldsymbol{\Sigma}_{r}},\hskip 5.0pt\mathbf{B}=\sqrt{\boldsymbol{\Sigma}_{r}}\mathbf{V}_{r}^{T},$
where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$. $\boldsymbol{\Sigma}_{r}$ is a diagonal matrix containing the largest $r$ singular values, and $\mathbf{U}_{r}$, $\mathbf{V}_{r}^{T}$ are corresponding singular vectors truncated from $\mathbf{U}$ and $\mathbf{V}^{T}$. This truncation and subsequent matrix formation let us approximate weight matrix $\mathbf{W}$ with two low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, thereby reducing the storage by $\frac{mr+rn}{mn}$.
@@ 3.1 Compressing the KV-Cache via Low-Rank Projection
As noted, a naive approach to compress the KV-Cache with low-rank projection is to apply SVD directly on the KV-Cache and store the top-$r$ singular vectors. However, this approach poses significant computational challenges during runtime that make it impractical for deployments (see Section 1).
To apply low-rank projection more efficiently, Palu uses SVD to decompose the key and value projection weight matrices. This approach is based on the observation that low-rank decomposition rewrites the linear projection layer from $\mathbf{y}=\mathbf{x}\mathbf{W}$ into $\mathbf{y}=\mathbf{x}\mathbf{A}\mathbf{B}$. Here, $\mathbf{A}\in\mathbb{R}^{d\times r}$ is the low-rank projection matrix, and $\mathbf{B}\in\mathbb{R}^{r\times d}$ is the reconstruction matrix derived by SVD. The forward process first down-projects the input token $\mathbf{x}\in\mathbb{R}^{d}$ into a low-dimensional latent space $\mathbf{h}\in\mathbb{R}^{r}$ and then up-projects it back to the original space:
@@
This two-step process lets Palu (1) store the lower dimension latent representation instead of the origin keys and values states, and (2) reconstruct them during decoding.
@@ 3.1.1 Integration with the Attention Mechanism and Offline Matrix Fusion
Based on the formula of attention output in equation 2, Palu absorbs the reconstruction matrix $\mathbf{B}^{v}_{i}$ into the output projection matrix $\mathbf{W}_{i}^{o}$ offline:
$\mathbf{a}_{i}\mathbf{W}^{o}_{i}$ $=(\mathbf{p}_{i}\mathbf{V}_{i})\mathbf{W}^{o}_{i}$ (5)
$=(\mathbf{p}_{i}\mathbf{H}^{v}_{i}\mathbf{B}^{v}_{i})\mathbf{W}^{o}_{i}$
$=\mathbf{p}_{i}\mathbf{H}^{v}_{i}(\mathbf{B}^{v}_{i}\mathbf{W}^{o}_{i})$
@@
$\mathbf{q}_{i}\mathbf{K}_{i}^{T}$ $=\mathbf{q}_{i}(\mathbf{H}^{k}_{i}\mathbf{B}^{k}_{i})^{T}$ (6)
$=\mathbf{x}_{t}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}(\mathbf{H}^{k}_{i})^{T}$
$=\mathbf{x}_{t}\Bigl{(}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}\Bigr{)}(\mathbf{H}^{k}_{i})^{T}.$
@@
3.1.2 Mechanism with Positional Embedding
Recent LLMs, such as the Llama family, apply positional embedding (i.e., RoPE (Su et al., 2021)) onto the query and key states before multiplication. The non-linear operations of adding positional embedding prevent the fusion for the attention score, as described in equation 6. In this case, Palu reconstructs the keys from latent representations on-the-fly during decoding. We improve reconstruction efficiency by designing an optimized kernel, which we present in Sec. 3.5.
Note that for some positional embedding methods, such as ALiBi Press et al. (2022), or new attention mechanisms, such as MLA DeepSeek-AI et al. (2024), positional embedding is not directly applied on the key states; therefore, the fusion of equation 6 remains valid. We expect the end-to-end speedup to be larger on these models. Fig. 2 presents a detailed workflow of Palu based on Llama’s attention with RoPE.
@@ 3.2.1 Multi-Head Low-Rank Decomposition
We name the per-head decomposition scheme in Sec.3.1.1 as multi-head low-rank decomposition (M-LRD). We found M-LRD often causes a non-negligible accuracy degradation (discussed further in Sec.4.1), possibly because SVD fails to capture the common information shared across heads. Therefore, alternative approaches are needed for preserving model accuracy.
@@ 3.2.2 Joint-Head Low-Rank Decomposition
Despite better preserving model information, J-LRD introduces significant computational and memory overhead during decoding. Specifically, the total number of floating point operations (FLOPs) to reconstruct the key or values state of one head now becomes $r_{\text{joint}}\cdot d_{h}\cdot n$. Assuming the same size as the total low-rank latent representations (i.e., $r_{\text{joint}}=\sum_{i=1}^{n}{r_{i}}$), the total reconstruction cost is $n$ times higher than M-LRD, whose total FLOPs is $r_{i}\cdot d_{h}\cdot n$. When considering the matrix fusion in Sec.3.1.1, the fused matrix size of J-LRD has a size of $r_{joint}\cdot d\cdot n$, which is also $n$ times larger than M-LRD, leading to a substantial higher memory consumption.

### v2

# Palu: KV-Cache Compression with Low-Rank Projection

@@ Abstract
Post-training KV-Cache compression methods typically either sample a subset of effectual tokens or quantize the data into lower numerical bit width. However, these methods cannot exploit redundancy in the hidden dimension of the KV tensors. This paper presents a hidden dimension compression approach called Palu, a KV-Cache compression framework that utilizes low-rank projection to reduce inference-time LLM memory usage. Palu decomposes the linear layers into low-rank matrices, caches compressed intermediate states, and reconstructs the full keys and values on the fly. To improve accuracy, compression rate, and efficiency, Palu further encompasses (1) a medium-grained low-rank decomposition scheme, (2) an efficient rank search algorithm, (3) low-rank-aware quantization compatibility enhancements, and (4) optimized GPU kernels with operators fusion. Extensive experiments with popular LLMs show that Palu compresses KV-Cache by 50%, while maintaining strong accuracy and delivering up to 1.89$\times$ speedup on the RoPE-based attention module. When combined with quantization, Palu’s inherent quantization-friendly design yields small to negligible extra accuracy degradation, while saving additional memory than quantization-only methods and achieving up to 2.91$\times$ speedup for the RoPE-based attention. Moreover, it maintains comparable or even better accuracy (up to 1.19 lower perplexity) compared to quantization-only methods. These results demonstrate Palu’s superior capability to effectively address the efficiency and memory challenges of LLM inference posed by KV-Cache. Our code is publicly available at: https://github.com/shadowpa0327/Palu
$\ast$\$$\ast$\$footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Correspondence to: Chi-Chih Chang, cc2869@cornell.edu
@@ 1 Introduction
Large language models (LLMs) have propelled AI into new applications and capabilities, providing a high-level intelligence that previous machine learning (ML) models could not achieve. To speed up inference, caching the Key-Value states (KV-Cache) in memory is a simple yet effective technique. However, the size of the KV-Cache can grow rapidly, straining memory capacity and bandwidth especially with long context lengths (Fu, 2024); further, the memory-bounded nature of the decoding stage limits inference speed when loading KV-Cache data (Gholami et al., 2024). Therefore, KV-Cache compression has become a central research topic for running LLMs efficiently.
Although emerging attention mechanisms such as Multi-Query Attention (MQA) (Shazeer, 2019), Group-Query Attention (GQA) (Ainslie et al., 2023) and Multi-head Latent Attention (MLA) (DeepSeek-AI et al., 2024) can reduce KV-Cache size, it either requires model pre-training or has a significant impact on model’s accuracy when converting from traditional Multi-Head Attention (MHA) (Chen et al., 2024). In contrast, post-training KV-Cache compression techniques offer an alternative approach to advance efficiency for existing models. Among various KV-Cache compression methods, quantization (Liu et al., 2024b; Hooper et al., 2024) and token eviction (Zhang et al., 2024; Xiao et al., 2024) stand out as effective strategies to reduce the memory footprint of KV-Cache.
Quantization methods aim to reduce the bit-width used to represent each piece of data, while token eviction techniques focus on retaining a partial set of KV-Cache. However, both methods neglect the hidden dimensions of the KV-Cache, where substantial redundancy often resides. To capitalize on this untapped potential, we introduce Palu, a post-training KV-Cache compression framework that leverages low-rank projection to reduce the hidden dimension of KV tensors, offering an additional and orthogonal compression dimension to existing quantization and token eviction methods.
@@
A naive way to utilize low-rank projection for compressing the KV-Cache is by directly mapping cached matrices into low-rank space (Jolliffe & Cadima, 2016; Zhao et al., 2024). However, this approach imposes an unacceptably heavy overhead of computing the decomposition matrices during runtime. To avoid this, Palu statically decomposes the Key and Value-projection weight matrices and caches the latent representations of the low-rank decomposition (see Fig. 1). This innovative design enables Palu to reduce memory while mitigating the runtime overhead of KV-Cache low-rank decomposition.
Figure 1: Palu’s low-rank projection method for KV-Cache reduction. A weight matrix W of linear projection is decomposed into two low-rank matrices. Input X is down-projected to latent representation H, which is cached. Y can be reconstructed from H using the up-projection matrix B.
In designing an effective decomposition strategy for attention modules with multiple attention heads, we observed a clear trade-off between accuracy and reconstruction overhead. Decomposing the projection matrices across all attention heads together improves accuracy by preserving global information, but this approach significantly increases reconstruction costs. On the other hand, decomposing each head separately reduces reconstruction overhead but leads to a higher loss in accuracy. To address this, Palu introduces a medium-grained, group-head low-rank decomposition that strikes a balance between accuracy and reconstruction efficiency.
@@
For LLMs, each linear projection module has a different sensitivity to compression (Sharma et al., 2023; Yuan et al., 2023). To exploit the sensitivity and improve accuracy, we design an efficient rank search algorithm based on Fisher information (Ly et al., 2017; Liu et al., 2021). Our algorithm automatically assigns a higher rank for important matrices and lower ranks for less critical ones, boosting accuracy at the same overall KV-Cache compression rate.
In addition to its low-rank decomposition, Palu is compatible with quantization techniques. We found that low-rank decomposition can introduce severe outliers in the latent representation, which significantly hinders accurate low-bit quantization. Although the Hadamard transformation has been shown to be effective for outlier elimination in recent studies (Tseng et al., 2024; Ashkboos et al., 2024b; Liu et al., 2024a; Chiang et al., 2024), its integration often introduces computational overhead during runtime. However, Palu’s inherent matrix pair structure makes it highly compatible with this technique, allowing the transformation matrices to be seamlessly fused into the forward and backward matrices, effectively mitigating outliers without impacting runtime efficiency.
@@
We evaluate Palu on widely used LLMs and benchmarks. Our experiments demonstrate that Palu maintains strong zero-shot accuracy and perplexity with up to 50% low-rank compression. Moreover, when combining low-rank compression with quantization, Palu achieves an impressive over 91.25% compression (11.4$\times$ reduction) and yields a significantly lower perplexity of 1.19 than KVQuant (Hooper et al., 2024), a state-of-the-art KV-Cache quantization method, which only achieves an 87.5% compression rate.
For latency evaluation, under a 50% KV-Cache compression rate without quantization, Palu demonstrates up to 1.89$\times$ and 2.2$\times$ speedup for RoPE-based and non-RoPE attention modules. When integrated with quantization, Palu achieves up to 2.91$\times$ and 6.17$\times$ acceleration on RoPE-based and non-RoPE attention, respectively. These results underscore Palu’s ability to significantly reduce KV-Cache memory footprint while boosting inference efficiency for LLMs.
@@
  • Palu, a new post-training KV-Cache compression framework that caches low-rank latent representations of Key and Value states.
  • Group-head low-rank decomposition (G-LRD), an optimization for balancing accuracy and reconstruction efficiency.
@@
  • A co-designed quantization compatibility optimization that eliminates low-rank-induced outliers and imposes zero runtime overhead.
@@ 2.2 Singular Value Decomposition (SVD)
SVD (Jolliffe & Cadima, 2016) is a commonly used technique for computing the low-rank approximation for a given matrix. We now introduce our specific use of SVD for Palu’s default low-rank decomposition method.
Given a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD decomposes $\mathbf{W}$ into three matrices: $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{T}$. Here, $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices containing the left and right singular vectors, respectively. The matrix $\bm{\Sigma}$ is a diagonal matrix that consists of singular values. We describe the decomposition as
$\mathbf{W}\approx\mathbf{A}\mathbf{B},\hskip 5.0pt\mathbf{A}=\mathbf{U}_{r}\sqrt{\bm{\Sigma}_{r}},\hskip 5.0pt\mathbf{B}=\sqrt{\bm{\Sigma}_{r}}\mathbf{V}_{r}^{T},$
where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$. $\bm{\Sigma}_{r}$ is a diagonal matrix containing the largest $r$ singular values, and $\mathbf{U}_{r}$, $\mathbf{V}_{r}^{T}$ are corresponding singular vectors truncated from $\mathbf{U}$ and $\mathbf{V}^{T}$. This truncation and subsequent matrix formation let us approximate weight matrix $\mathbf{W}$ with two low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, thereby reducing the storage by $\frac{mr+rn}{mn}$.
@@ 3.1 Compressing the KV-Cache via Low-Rank Projection
To apply low-rank projection more efficiently than directly decomposing the KV-Cache during runtime, Palu uses SVD to decompose the Key and Value projection matrices. This approach is based on the observation that low-rank decomposition rewrites the linear projection layer from $\mathbf{y}=\mathbf{x}\mathbf{W}$ into $\mathbf{y}=\mathbf{x}\mathbf{A}\mathbf{B}$. Here, $\mathbf{A}\in\mathbb{R}^{d\times r}$ is the low-rank projection matrix, and $\mathbf{B}\in\mathbb{R}^{r\times d}$ is the reconstruction matrix derived by SVD. The forward process first down-projects the input token $\mathbf{x}\in\mathbb{R}^{d}$ into a low-dimensional latent space $\mathbf{h}\in\mathbb{R}^{r}$ and then up-projects it back to the original space:
@@
This two-step process lets Palu (1) store the lower dimension latent representation instead of the origin key and value states, and (2) reconstruct them during decoding.
@@ 3.1.1 Integration with the Attention Mechanism and Offline Matrix Fusion
Based on the formula of attention output in Eq. 2, Palu absorbs the reconstruction matrix $\mathbf{B}^{v}_{i}$ into the output projection matrix $\mathbf{W}_{i}^{o}$ offline:
$\mathbf{a}_{i}\mathbf{W}^{o}_{i}$ $=(\mathbf{p}_{i}\mathbf{V}_{i})\mathbf{W}^{o}_{i}=(\mathbf{p}_{i}\mathbf{H}^{v}_{i}\mathbf{B}^{v}_{i})\mathbf{W}^{o}_{i}=\mathbf{p}_{i}\mathbf{H}^{v}_{i}(\mathbf{B}^{v}_{i}\mathbf{W}^{o}_{i})$ (5)
@@
$\mathbf{q}_{i}\mathbf{K}_{i}^{T}$ $=\mathbf{q}_{i}(\mathbf{H}^{k}_{i}\mathbf{B}^{k}_{i})^{T}=\mathbf{x}_{t}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}(\mathbf{H}^{k}_{i})^{T}=\mathbf{x}_{t}\Bigl{(}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}\Bigr{)}(\mathbf{H}^{k}_{i})^{T}.$ (6)
@@
3.1.2 Compability with Positional Embedding
Recent LLMs, such as the Llama family, apply Rotary Positional Embedding (i.e., RoPE (Su et al., 2021)) onto the Query and Key states prior to their multiplication. The non-linear nature of these positional embeddings prevents the matrix fusion of attention scores, as outlined in Eq. 6. To address this, Palu dynamically reconstructs the keys from latent representations during decoding. We enhance reconstruction efficiency with a custom-designed GPU kernel, detailed in Sec. 4.1 and further described in Appendix B.
Note that for some positional embedding methods, such as ALiBi (Press et al., 2022), or new attention mechanisms, such as MLA (DeepSeek-AI et al., 2024), positional embedding is not directly applied to the Key states. Consequently, the fusion described in Eq. 6 remains valid. For these non-RoPE attention modules, Palu achieves greater speedup compared to RoPE-based attention, as their reconstruction can be avoided with matrix fusion. A detailed workflow of Palu in the context of Llama’s RoPE-based attention is illustrated in Fig. 2.
@@ 3.2.1 Multi-Head Low-Rank Decomposition
We name the per-head decomposition scheme in Sec. 3.1.1 as multi-head low-rank decomposition (M-LRD). We found M-LRD often causes a non-negligible accuracy degradation (discussed further in Sec. 4.2), possibly because SVD fails to capture the common information shared across heads. Therefore, alternative approaches are needed to preserve model accuracy.
@@ 3.2.2 Joint-Head Low-Rank Decomposition
Despite better preserving model accuracy, J-LRD introduces significant computational and memory overhead during decoding. Specifically, the total number of floating point operations (FLOPs) to reconstruct the Key or Value state of one head now becomes $r_{\text{joint}}\cdot d_{h}\cdot n$. Assuming the same size as the total low-rank latent representations (i.e., $r_{\text{joint}}=\sum_{i=1}^{n}{r_{i}}$), the total reconstruction cost is $n$ times higher than M-LRD, whose total FLOPs is $r_{i}\cdot d_{h}\cdot n$. When considering the matrix fusion in Sec. 3.1.1, the fused matrix of J-LRD has a size of $r_{joint}\cdot d\cdot n$, which is also $n$ times larger than M-LRD, leading to substantial higher memory consumption.